OpenStudy (anonymous):

6. Does the following equation represent growth or decay? (Points : 1)

OpenStudy (anonymous):

If the exponent is negative... decay. If positive... growth.

OpenStudy (anonymous):

Not necessarily.... if 0<b<1, say b=1/2, then \(\large y=(\frac{1}{2})^{-x}=(2^{-1})^{-x}=2^x \) is a growth equation.